OpenStudy (anonymous):

add (1/8x^4 + 5/6x^3 +3/8 +8) + (-5/8x^4 +5/8x^2 -8)

OpenStudy (anonymous):

\[(1/8x^4 + 5/6x^3 +3/8 +8) + (-5/8x^4 +5/8x^2 -8)\] \[=-4/8x^4+5/6x^3+5/8x^2+3/8\] \[=-1/2x^4+5/6x^3+5/8x^2+3/8\]

OpenStudy (anonymous):

I show that is the wrong answer :(